OpenStudy (anonymous):

Find the absolute minimum value of f(x) =(3/2)x + 2/(9x^2) on (0 ,∞).

hero (hero):

I suppose graphing it is out of the question

hero (hero):

I found the minimum value. Are you there?

OpenStudy (anonymous):

i try taking the derivative then i set that equal to zero but then i can never get past that

hero (hero):

f(x) = (3/2)x + 2/(9x^2) f'(x) = 3/2 - 4/(9x^3) 0 = 3/2 - 4/(9x^3) 4/(9x^3) = 3/2 Cross multiply: 2(4) = 3(9x^3) 8 = 27x^3 8/27 = x^3 cube root(8/27) = x 2/3 = x Plug x = 2/3 back into f(x): f(2/3) = (3/2)(2/3) + 2/(9(2/3)^2) f(2/3) = 1 + 2/(9(4/9)) f(2/3) = 1 + 2/4 f(2/3) = 1 + 1/2 f(2/3) = 2/2 + 1/2 f(2/3) = 3/2 Thus the minimum value of f(x) is 3/2

hero (hero):

If you have any questions about any of that, let me know

OpenStudy (anonymous):

OHHHHH, ok i had gotten to cube root(8/27) then thats where i got stuck! That makes sense, Thankyou so much! :]